### v2

# ConceptWM: Watermarking Diffusion Models for Visual Concept Copyright Protection

Abstract.
The personalization techniques of diffusion models succeed in generating specific concepts but also pose threats to copyright protection and illegal use. Model Watermarking is an effective method to prevent the unauthorized use of subject-driven or style-driven image generation, safeguarding concept copyrights. However, under the goal of concept-oriented protection, current watermarking schemes typically add watermarks to all images rather than applying them in a refined manner targeted at specific concepts. Additionally, the personalization techniques of diffusion models can easily remove watermarks. Existing watermarking methods struggle to achieve fine-grained watermark embedding with a few images of specific concept and prevent removal of watermarks through personalized fine-tuning. Therefore, we introduce a novel concept-oriented watermarking framework that seamlessly embeds imperceptible watermarks into the concept of diffusion models. We introduce Fidelity-preserving Latent Watermarking (FLW) to generate latent watermarks based on image characteristics and the Adversarial Watermarking Modulation module to prevent ”jailbreaking” via personalized finetuning. To enhance U-Net’s efficiency in learning watermark patterns with limited samples, we propose Efficient Concept Watermark Finetuning, which alternates optimization of model parameters for both watermark embedding and concept learning. We conduct extensive experiments and ablation studies to verify our framework. Our code is available at https://anonymous.4open.science/r/Conceptwm-4EB3/.
Latent Diffusion Model, Image Generation, Model Watermarking
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer vision
1. Introduction
Recent advancements in diffusion models have significantly upgraded image generation techniques, enabling the creation of realistic images from simple text descriptions and meeting the demand for personalized and customized content. The upgraded diffusion models excel in generating novel concept descriptions, such as unique visual themes, motifs, stylistic preferences, or characters, using only a few images of specific themes or styles (Shi et al., 2024; Gal et al., 2023, 2022; Kumari et al., 2023). Through personalization tools, e.g., DreamBooth (Ruiz et al., 2023) and Lora (Hu et al., 2021), any users with minimal design skills can create images of specific concepts by using only a few personal images and simple text inputs. However, the use of personalization techniques introduces potential risks of concept misuse or leakage. Untrustworthy users may exploit these tools to create unauthorized concepts, leading to violations of original creators’ intellectual property and creativity. For instance, creating concepts related to celebrities or artworks can result in deepfake incidents (Juefei-Xu et al., 2022), the spread of fake news and online spam (Barrett et al., 2023), and serious copyright violations.
Figure 1. (a) Post-processing Watermark. (b) U-net Finetuning Watermark. (c) VAE-Finetuning Watermark. (d) Latent-based Watermark. (e) Ours.
Concept watermarking technique is considered to be an alternative to address ownership and accountability issues of specific concepts in generative model, which aims to embed watermark information into the concept and extract the watermarks from subsequent generated images on the basis of the watermarked concept (Feng et al., 2023). In general, the model owner possesses a pre-trained diffusion model and aims to generate images for a new concept. The concept owner provides images of the specific concept to fine-tune the existing diffusion model. During this fine-tuning process, watermark information related to the concept is embedded, enabling the model to generate images with the new concept. A watermark decoder can extract the watermark information from the generated images.
However, as shown in Fig. 1, existing watermarking methods have not yet to achieve concept watermarking, despite the growing interest in adding watermarks to generated images (Kim et al., 2024; Xiong et al., 2023; Zhang et al., 2024; Yang et al., 2024; Fernandez et al., 2023; Zhao et al., 2023b; Wen et al., 2024; Feng et al., 2024; Cox et al., 2007). On the one hand, prior work has rarely addressed fine-grained watermarking mechanisms for specific concepts. Most existing methods (Wen et al., 2024; Fernandez et al., 2023; Zhao et al., 2023b) watermark the entire images rather than images of specific concepts, making them inapplicable for concept watermarking scenario. A major challenge in concept watermarking is to ensure that the watermark not only exists within a specific image but also remains present across various image variations genraeted by the diffusion model. Compared to traditional image watermarks embedded at the pixel level, concept watermarks appears a higher-level technical barrier, as it associates the watermarks with high-level sematic features related to specific concepts. Since the number of images with a specific concept is limited, embedding watermarks in personalization models is a great challenge, especially when fine-tuning the model by injecting a new concept.
On the other hand, robustness remains a challenge for current watermarking methods in concept protection scenarios, when considering diverse adversaries. Recent studies have demonstrated that existing methods are ineffective against model purification attacks (Zhao et al., 2024; Liu et al., 2024), as adversaries can bypass watermarking schemes by generating watermark-free images through personalized fine-tuning. The technical challenge lies in balancing adversariality, image quality, and watermark fidelity. One potential approach to enhancing watermark adversariality is to inject imperceptible protective adversarial perturbations into images, even though this type of methods frequently compromises both image quality and watermark fidelity (Liang et al., 2023; Shan et al., 2023; Van Le et al., 2023). Therefore, there is an urgent need for watermarking solution that ensures concept tracing while effectively addressing adversarial threats.
Our investigation finds three fundamental properties required by model watermarking, including refined nature, robustness, and adversarial properties. To be specific, refined nature, as a core component of diffusion model, requires the ability to embed a watermark into a specific concept through minimal training data and steps without compromising fidelity, as the knowledge of a specific concept generally is learnt by the denoising network (U-Net). Robustness ensures that the watermark remains stable across various image processing operations and generation parameters. Adversarial properties typically guarantee that the generated concept-specific images cannot be easily “jailbreaken” via personalized fine-tuning.
To address these challenges, we propose a novel concept watermarking framework, named ConceptWM, which focuses on two critical factors, namely refined watermark training and adversarial perturbation modulation. To achieve the concept watermark embedding, we develop a Efficient Concept Watermark Finetuning (ECWF) scheme, which rapidly learns watermark embedding patterns in diffusion models with limited data and training resources. Additionally, we develop a Fidelity-preserving Latent Watermark (FLW) module that incorporates the watermark into the latent space and employs a distortion layer in order to enhance robustness against image processing attacks. Next, we propose an Adversarial Watermark Modulation (AWM) to balance adversarial perturbation, watermark fidelity, and image quality, while mitigating the risk of adversarial attacks bypassing the watermark through personalized fine-tuning. This scheme utilizes an alternating optimization strategy to refine both referenced model parameters and adversarial perturbations, so that an optimal equilibrium among adversarial robustness, image quality, and watermark fidelity can be achieved.
Figure 2. Framework of the proposed ConceptWM watermarking scheme for Latent Diffusion Model. The Efficient Concept Watermark Finetuning module employs a two-step training process to learn specific concepts and watermark embedding patterns. The Adversarial Watermark Modulation module utilizes a surrogate model fully fine-tuned on a referenced clean image set $X_{B}$ to guide PGD optimization, iteratively generating adversarial watermarks through multiple rounds of watermark and noise integration.
Our contributions are as follows: (1) We propose a novel concept-driven watermarking framework that seamlessly embeds imperceptible watermarks into the concept of diffusion models. The proposed framework highlights the importance of refined watermark training and fidelity-preserving perturbations in real-world applications of diffusion model watermarking. (2) We propose a mechanism named Adversarial Watermark Modulation that endows concept watermarks with adversarial properties. Through the alternating optimization of model training and adversarial algorithms, it effectively balances adversarial robustness, image quality, and watermark fidelity. (3) Experiments demonstrate that ConceptWM effectively integrates adversarial perturbations into watermarks and maintains better robustness and accuracy in concept watermark embedding and extraction compared with existing 4 baselines over 6.3% image processing attacks.
2. Related Work
Personalization Techniques for Diffusion Models. Recent explorations mainly focus on injecting specific concepts into the model through fine-tunning tools, e.g., Textual Inversion (Gal et al., 2022), DreamBooth (Ruiz et al., 2023) and LoRA (Low-Rank Adaptation)(Hu et al., 2021), given the substaintial computational requirements of training Stable Diffusion from scratch. Specifically, Textual Inversion focuses on training text embedding so that concepts are allowed to be injected into the text encoder without altering weights of the U-Net. Dissimilar to conventioanl text-to-image fine-tuning, DreamBooth fine-tunes the entire U-Net portion of the Stable Diffusion model and employs prior loss during training to prevent overfitting. LoRA is designed to achieve lightweigth efficient fine-tuning by training incremental weights in the attention layers of the U-Net. Our work focuses on watermarking specific concepts generated through DreamBooth fine-tuning.
Model Watermarking for Latent Diffusion Models. The embedding characteristics of different watermarking methods for latent diffusion models are illustrated in Figure 1. It is mainly divided into Post-processing Watermark, U-Net Finetuning Watermark, VAE-Finetuning Watermark, and Latent-based Watermark. These watermarking methods aim to protect the copyright of images or generative models, so they often embed watermarks into all images. For example, WDM (Zhao et al., 2023b) is designed to train an autoencoder to stamp a watermark on all training data before re-training the generator from scratch . Aqualora (Feng et al., 2024) is a lora-based method that watermarks all images. VAE-based methods (Xiong et al., 2023; Kim et al., 2024; Fernandez et al., 2023) fine-tune VAE-Decoder to ensure all generated images containing the watermark. However, this all-image type of methods encounters challenges in protecting specific styles. A large volume of images are generally required for the model to learn the watermark embedding pattern, but concept-oriented images for finetuning are limited to only a few.
Protective Perturbation against Stable Diffusion. Recent studies mostly have addressed disrupting the fine-tuning process by adding imperceptible protective noise to images, in order to prevent personal images from potential infringement during the fine-tuning of Stable Diffusion. For example, AdvDM employs adversarial noise to safeguard personal images on Stable Diffusion by maximizing the Mean Squared Error loss during optimization (Liang et al., 2023). Anti-DreamBooth (Van Le et al., 2023) offers a bi-level min-max optimization for generating protective perturbations, through integrating the DreamBooth fine-tuning process of Stable Diffusion into its framework. Other studies (Wu et al., 2023; Ye et al., 2023; Zhao et al., 2023a; Zheng et al., 2023) have explored similar adversarial perturbation techniques to generate protective noise for images.
3. Methodology
3.1. Overview
The main objective of our scheme is to protect images of a specific concept that are generated by the diffusion model, by embedding an adversarial watermark pattern into the U-Net component of the diffusion model. Fig. 2 illustrates an overview of our approach. To embed a watermark that is both fidelity-preserving and conducive to learning by U-Net, we develop a Fidelity-preserving Latent Watermarking (FLW) module, which adds a combination layer for the message encoder and latent concatenation, so that watermarks can be generated based on image characteristics rather than a fixed watermark signal. To protect the generated concept-specific images from being “jailbroken” through personalized finetuning, we develop an Adversarial Watermarking Modulation module to integrate adversarial propeties into the watermark. To enable U-Net to efficiently learn watermark patterns during concept learning with limited samples, we present an Efficient Concept Watermark Finetuning module that alternates optimization of model parameters to simultaneously achieve adversarial watermark embedding and concept learning.
3.2. Fidelity-preserving Latent Watermark
Figure 3. The comparison of two variants of Latent watermark namely Overlay-based Latent Watermarking and Fusion-based Latent Watermarking.
Existing image watermarking techniques frequently arise substantial discrepancies between the watermark embedding pattern and the latent space. This mismatch issue may cause the disruption or even complete loss of watermark information when the image is encoded into the latent space by a VAE encoder or regenerated through diffusion, as has been discussed by recent watermarking attacks (An et al., 2024; Ballé et al., 2018; Cheng et al., 2020). Thus, we aim to embed a robust watermark within the image that is both resilient and conducive to efficient learning for U-Net. To be specific, we have developed two variants of latent watermarking, which are Overlay-based Latent Watermarking (OLW) and Fidelity-preserving Latent Watermarking (FLW). As shown in Fig. 3, both variants involve the latent domain perturbations. The watermarked images are processed through the message decoder to evaluate the discrepancy between the decoded watermark information and the original content of the watermarked image. The message loss is the Binary Cross Entropy (BCE) between the message $m$ and $\sigma(m^{\prime})$. Given the message length $N$, the message loss $\mathcal{L}_{m}$ is defined in Eq. (1).
(1) $\mathcal{L}_{m}=-\sum_{k=1}^{N}m_{k}\log\sigma(m_{k}^{\prime})+(1-m_{k})\log(1-\sigma(m_{k}^{\prime}))$
Overlay-based Latent Watermarking. OLW essentially overlays a fixed watermark signal that can be removed through subtraction. We observe that OLW-watermarked images may face the issue of the watermark integrity, since noticeable artifacts appear in local regions of the images, making the images distinguishable to the huaman eye.
Fidelity-preserving Latent Watermarking. To address issues of OLW, we have developed Fidelity-preserving Latent Watermarking (FLW) that adds a combination layer after the message encoder and latent concatenation. To ensure visual consistency, we calculate LPIPS loss $\mathcal{L}_{pips}$ (Zhang et al., 2018) and utilize a Peak Pixel Difference (PPD) Loss $\mathcal{L}_{ppd}$ which selects the most prominent parts of the image and calculates channel loss to reduce excessive modification of the images. Our training objective ($\mathcal{L}_{total}$) can be expressed in Eq. (2), where $\lambda$ and $\mu$ represent the coefficients.
(2) $\mathcal{L}_{total}=\mathcal{L}_{m}+\lambda\mathcal{L}_{pips}+\mu\mathcal{L}_{ppd}$
The jointly optimized message encoder and decoder, working in concert with the Adversarial Watermark module, will empower the U-Net to learn robust watermarking patterns.
3.3. Adversarial Watermark Modulation
To protect the generated concept-specific images from “jailbreaking” through personalized fine-tuning, we integrate adversarial properties into the watermark. By maximizing the image generation loss with adversarial noise added to the watermark and minimizing the concept learning loss through Dreambooth fine-tuning with reference images, we ensure that the adversarial watermark does not interfere with concept learning while preventing “jailbreaking” through personalized fine-tuning.
Image Generation Loss. To enable the conditional denoising model $\epsilon_{\theta}(x,t,c)$ to match the data distribution of the target images, $\mathcal{L}_{cond}$ is a loss function that ensures the conditional denoising model generates images conditioned on a given text condition ($c$). Given the input image $x_{0}$, the $\mathcal{L}_{cond}$ is as follows in Eq. (3):
$\mathcal{L}_{cond}(\theta,x_{0})=\mathbb{E}_{x_{0},t,c,\epsilon\in\mathcal{N}(0,1)}\|\epsilon-\epsilon_{\theta}(x_{t+1},t,c)\|_{2}^{2}$
where each variable $x_{t}$ is constructed through noise injection at a corresponding timestep $t$ within the set $\{1,2,\cdots,T\}$.
[…]
$\mathcal{L}_{ft}(\theta,x_{0})=\mathbb{E}_{x_{0},t}\|\epsilon-\epsilon_{\theta}(x_{t+1},t,c)\|_{2}^{2}+\lambda\|\epsilon^{\prime}-\epsilon_{\theta}(x_{t+1}^{\prime},t,c_{cls})\|_{2}^{2}$
where $\epsilon$ and $\epsilon^{\prime}$ are both sampled from the standard normal distribution $\mathcal{N}(0,\mathbf{I})$, The noisy variable $x_{t+1}^{\prime}$ is generated from the original stable diffusion model based on the prior prompt $c_{\text{cls}}$. $\lambda$ emphasizes the importance of the prior term in the generation process. The term $\lambda$ emphasizes the importance of the prior term in the generation process.
Algorithm 1 Adversarial Watermark Modulation
1:experiment AWM(($x_{0}$,$X_{B}$,$\theta$))
2:     for  $x_{b}\in X_{B}$ do
3:         c= “ a photo of $\langle$ concept $\rangle$ person ”
4:         $\theta=\operatorname{TrainOneStepDreamBooth}(\theta,x_{b},c)$
5:     end for
6:     $s\leftarrow secret$
7:     $x_{w}=LatentWatermarkEmbedding(x_{0},s)$
8:     $\sigma_{1}=x_{w}-x_{0}$
9:     $Initialize$ $x_{0}^{0}\leftarrow x_{0}$
10:     for $t=1\ldots T$ do
11:         $\sigma_{2}^{*(i)}=\alpha\cdot sign(\nabla_{x}\mathcal{L}_{cond}(\theta,x_{0}^{t}+\sigma_{1}))$
12:         ${x_{0}}^{t+1}={x_{0}}^{t}+\sigma_{2}^{*(i)}$
13:     end for
14:     $\varphi=x_{0}^{T}+\sigma_{1}-x_{0}$
15:return $\varphi$
Adversarial Watermark Modulation. The goal of this modulation is to maximize $L_{cond}$ in Eq. (3) to hinder personalized finetuning and minimize $L_{ft}$ in Eq. (4) to learn the concept. Our solution is to use watermark-free images $\mathcal{X}_{A}$ for one-step DreamBooth training to reduce the DreamBooth loss $\mathcal{L}_{\text{ft}}$, while simultaneously applying Projected Gradient Descent (PGD) (Madry, 2017) to enhance the adversarial loss $L_{cond}$. The model used for PGD is trained with clean referenced images $\mathcal{X}_{B}$ in a one-step dreambooth training process, rather than the images used for concept learning. The purpose of this approach is to allow the model to learn adversarial perturbations without affecting the learning of the concept. The specific procedure is defined by Eq.s (5) and (6), where $\sigma_{1}$ denotes the residual of the watermarked image and original image. $x^{i}$ denotes the clean samples.
(5) $\theta^{*}=\arg\min_{\theta}\sum_{x\in\mathcal{X}_{A}}\mathcal{L}_{ft}(\theta,x^{i})$
(6) $\varphi^{(i)}=\arg\max_{\sigma_{2}^{(i)}}\mathcal{L}_{cond}(\theta^{*},x^{(i)}+\sigma_{1}+\sigma_{2}^{(i)})$
The adversarial noise $\sigma^{*(i)}$ is within the constraint range of $\eta$. The Adversarial Watermark Modulation outputs an adversarial watermark $\varphi$ as the parameter for Concept-oriented Watermark Fine-tuning to facilitate learning.
Algorithm 2 Efficient Concept Watermark Finetuning
1:experiment $CWF$($X_{A}$,$X_{B}$,$\theta$,E)
2:     for  $x_{0}\in X_{A}$ do
3:         c=“ a photo of $\langle$ concept $\rangle$ person ”
4:         $\theta=\operatorname{TrainOneStepDreamBooth}(\theta,x_{0},c)$
5:         $\theta^{*}\leftarrow\text{ copy parameters from }\theta$
6:         $\varphi=AWM(x_{0},X_{B},\theta)$
7:         $t\sim\mathrm{Uniform}(\{1,\ldots,T\})$
8:         $z_{0}=E(x_{0})$
9:         $z_{t}\leftarrow\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$
10:         $z_{t}^{\mathrm{w}}\leftarrow\sqrt{\bar{\alpha}_{t}}(z_{0}+\varphi)+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$
11:         Take gradient step on $\nabla_{\theta}\|{\epsilon_{\theta}}(z_{t}^{\mathrm{w}},t,c)-{\epsilon_{\theta^{*}}}(z_{t},t,c)\|^{2}$
12:     end for
13:return $\theta$
[…]
3.4. Efficient Concept Watermark Finetuning
We aim to integrate adversarial watermarking into concept learning in this module. The primary challenges are as follows: The availability of fine-tuned images for specific concepts is often limited, hindering the model’s ability to learn the distinct watermark patterns. The introduction of adversarial perturbations can interfere with the DreamBooth training process, as these protective perturbations prevent the model from effectively learning the concept. Our core idea is to decouple concept learning from adversarial watermark learning. In the DreamBooth stage, the U-Net learns the specific concept. In the adversarial watermark learning stage, the model is fed with both watermarked and non-watermarked images, and by using the difference in the predictions as a loss, it guides the generative model in learning the watermark embedding pattern.
The training objective of this stage is defined by Eq. (7), where $\varphi$ represents the adversarial perturbation containing the watermark generated through AWM, $\epsilon_{\theta}^{*}$ denotes the fixed model, and $\epsilon_{\theta}$ refers to the trained model.
(7) $\mathcal{L}_{\mathrm{cwft}}(\theta,x_{t})$ $=\mathbb{E}_{t,c,x}\Big{[}\|\boldsymbol{\epsilon}_{\theta}(x_{t}+\varphi,t,c)-\boldsymbol{\epsilon}_{\theta}^{*}(x_{t},t,c\big{)}\|^{2}\Big{]}$
Algorithm 2 outlines the finetuning procedure for the U-Net denoising model $\theta$ with three components: target concept images $X_{A}$, clean reference images $X_{B}$, and the VAE encoder $E$. The process has three phases: In concept specialization phase (Lines 1-5), The model $\theta$ is fine-tuned on $X_{A}$ to produce a specialized model $\theta^{*}$ that captures the target concept while maintaining general denoising capabilities. In adversarial watermark generation phase (Line 6), an adversarial perturbation is computed through the Fidelity-preserving Latent Watermark in section 3.2. In watermark finetuning phase (Lines 7-13), the model is fed with both watermarked and non-watermarked images, and the difference in the predictions is used as a loss to guide the generative model in learning the watermark embedding pattern.
[…]
Figure 4. Demonstration of the generated images. Below each image, we display the corresponding Bit Accuracy, along with the image quality metrics PIQE and NIQE.
4. Experiment
4.1. Implementation Details
Datasets. To evaluate the effectiveness of the proposed methods, for subject-driven generation, we select the face dataset CelebA-HQ (Zhu et al., 2022). For style-driven generation, we select 12 artists with distinct styles from WikiArt  [^1] https://www.wikiart.org, with each artist contributing 10 artworks that are consistent in terms of both artistic style and time period. Each image is resized and center-cropped to a resolution of 512 × 512. In each experiment, a total of 5–8 carefully selected images were utilized for DreamBooth fine-tuning.
[…]
Type Method Concept Oriented Concept Protection Fidelity Robustness
FID $\uparrow$ CLIP $\uparrow$ BitAcc $\uparrow$ BitAcc(Adv.) $\uparrow$ TPR $\uparrow$ TPR (Adv.$\uparrow$
None - - 24.37 0.367 - - - -
Post-Processing DwtDctSvd $\times$ $\times$ 25.01 0.363 100.0 71.47 1.000 0.367
RivaGan $\times$ $\times$ 24.72 0.363 98.47 83.21 0.994 0.651
Latent-based Tree-Ring $\times$ $\times$ 25.69 0.364 - - 1.000 0.821
VAE-based StableSig. $\times$ $\times$ 24.92 0.362 98.49 81.94 0.993 0.848
Concept-Oriented Ours(OLW-based) 24.06 0.357 98.68 90.03 0.998 0.884
Ours(FLW-based) 24.58 0.365 95.68 89.32 0.994 0.858
Table 1. A comparison between our method and previous watermarking techniques is presented. The capacities for DWT-DCT-SVD, RivaGAN, StableSignature, and our method are 48 bits, 32 bits, 48 bits, and 48 bits, respectively. We control the false positive rate (FPR) at $10^{-5}$ and evaluate the true positive rate (TPR). Since Tree-ring is a zero-bit watermark, bit accuracy cannot be calculated in this case. ”Adv.” (Adversarial) refers to the average performance of images under various distortions.
Distortion Layer during training. During training, several data augmentation techniques were employed to enhance the model’s robustness. We follow the setup of attack layers in the representative work Stable Signature (Fernandez et al., 2023). Specifically, JPEG compression was simulated by using the HiDDeN (Zhu et al., 2018) model’s JPEG distortion layer, with compression quality randomly chosen between 40 and 100 to simulate different levels of image compression. Random cropping was applied, resizing the cropped image within the range of [256, 512] pixels, and then scaling it back to 512×512 pixels to ensure variability in the region of interest. Gaussian blur was applied with kernel sizes randomly selected between [3, 9] and blur intensities ranging from 0 to 2, simulating distortions from defocus or motion blur. Gaussian noise with a mean of 0 and variance of 10 was added to simulate low-quality or noisy images. Color jittering adjusted brightness, contrast, and saturation within the range of [0.8, 1.25], and hue was randomly shifted between [-0.2, 0.2] to simulate lighting and color variations. These augmentations effectively improved the model’s adaptability to compression, noise, blur, and color variations, ensuring stable watermarking even under diverse image transformations.
Evaluation Metrics. We evaluate the detection performance using the true positive rate (TPR) at a false positive rate (FPR) of 10e-5. For traceability performance, we assess bit accuracy. To measure the quality of image generation, we compute the Peak Signal-to-Noise Ratio (PSNR) (Hore and Ziou, 2010) and Structural Similarity Index (SSIM) (Wang et al., 2004) to evaluate image distortion; and the Natural Image Quality Evaluator (NIQE) (Mittal et al., 2012) and Perceptual Image Quality Evaluator (PIQE) (Venkatanath et al., 2015) for evaluating overall image quality.
Watermarking baselines. We select four typical baselines: two official watermark of Stable Diffusion (Rombach et al., 2022) called DwtDctSvd (Cox et al., 2007), and RivaGAN (Zhang et al., 2019),VAE-based method Stable Signature (Fernandez et al., 2023) and Latent-based method Tree-ring (Wen et al., 2024). DwtDctSvd, RivaGAN, and Stable Signature all use 48-bit watermarks, while Tree-ring uses P-value as the evaluation method.
Metrics Methods Distortions
Contrast Brightness Blur Noise Jpeg Crop&Rescale Vae.C Denoising Average
Bit Accuracy(%$\uparrow$ Dwtdctsvd 88.41 86.80 92.42 65.87 83.32 54.02 55.74 49.53 72.01
RivaGan 97.96 94.67 94.89 91.09 90.68 94.17 70.62 58.64 86.59
Stablesig. 97.86 95.48 89.88 78.94 84.40 97.47 69.94 52.58 83.31
Ours (FLW-based) 94.26 93.46 93.68 92.42 93.63 90.26 90.12 84.63 91.34
Tpr(Fpr=$10^{-5}$$\uparrow$ Dwtdctsvd 0.894 0.752 0.943 0.054 0.734 0.008 0.010 0.000 0.424
RivaGan 0.989 0.956 0.979 0.926 0.928 0.956 0.284 0.010 0.754
Stablesig. 0.987 0.985 0.905 0.369 0.821 0.989 0.293 0.000 0.669
Tree-ring 1.000 1.000 0.954 0.701 0.905 0.121 0.958 1.000 0.830
Ours (FLW-based) 0.968 0.951 0.947 0.958 0.949 0.912 0.910 0.734 0.916
Table 2. The performance of ConceptWM against various types of attacks is compared with other watermarking schemes. The attacks include (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al., 2020) and (8) Diffusion-based Reconstructive Attack (Zhao et al., 2023c).
4.2. Watermark Performance Comparison
We evaluate our method against 4 typical baselines on two tasks: (1) Detection: All methods are tested as single-bit watermarks with a unified watermark, with the False Positive Rate (FPR) is set to $10^{-5}$, and we assess the True Positive Rate (TPR) on 1,000 watermarked images. (2) Traceability: Except for the single-bit watermark Tree-Ring, all other methods are tested as multi-bit watermarks. Specifically, we simulate a scenario with 1,000 model users, each requiring a distinct watermark for model tracing. Each user generates 10 images, creating a dataset of 10,000 watermarked images. During testing, we calculate the Bit Accuracy by comparing the watermark of each user with the watermarked images generated by the user. The user with the highest Bit Accuracy is considered the traced user.
The comparison results, shown in Table 1, demonstrate that our watermarking method achieves superior robustness, significantly outperforming the baselines in both tasks. In terms of Bit Accuracy, our method outperforms the best baseline by around 6.3%. OLW exhibits strong performance in robustness experiments. However, it suffers from relatively unsatisfactory image quality (FID 24.06,CLIP 0.357). FLW demonstrates a relatively balanced result, maintaining good image quality while ensuring strong robustness. We found that there is a trade-off between image quality, robustness, and detection accuracy in watermark embedding. Due to extensive adversarial training on the watermark model, FLW achieves a bit accuracy of 95.68 in a non-attack environment, which is relatively lower compared to other watermarking schemes. To further improve bit accuracy in a non-attack setting, it would require considering either lower image quality or reduced robustness.
4.3. Fidelity
Table 1 presents a comparison of our method with other baseline methods. For fidelity evaluation metrics, we use the Fréchet Inception Distance (FID) (Heusel et al., 2017) for assessing image diversity and semantic relevance, which is calculated on 5000 images from the COCO 2017 validation set (Lin et al., 2014). Additionally, we use CLIP score (Radford et al., 2021), a method for measuring image similarity, which provides results more aligned with human judgment.
Settings BitAcc(%) TPR (Fpr=$10^{-6}$)
Steps 25 0.917 0.921
50 0.936 0.946
100 0.936 0.951
Sampler DDIM 0.929 0.939
DPM-M 0.938 0.972
EULER 0.922 0.941
HEUN 0.947 0.962
Guidance Scale 5 0.932 0.948
7.5 0.917 0.932
10 0.933 0.947
Generation Size 512 0.933 0.951
768 0.812 0.748
VAE V2.1 0.939 0.954
V1.4 0.901 0.913
Table 3. The impact of inference settings on ConceptWM watermark detection.
Main Results. The performance of this method outperforms other methods in metrics like CLIP and FID, with minimal performance difference compared to the original model (FID difference of 0.21, CLIP difference of 0.02). This method is a localized fine-tuning of the model and does not significantly impact the model’s overall performance. Due to the smaller dataset and the more refined watermarking method, our method has a smaller impact on model generation quality compared to other methods. It can also be observed that, among the two variants of latent watermarking introduced, OLW and FLW, the image quality produced by OLW, with an FID of 24.06 and semantic quality (CLIP) of 0.357, is comparatively lower. In contrast, the image quality produced by FLW, has an (FID) of 24.58 and semantic quality (CLIP) of 0.365. This disparity arises because the watermark in FLW is contextually related to the image content, whereas the watermark in OLW is a fixed, image-independent watermark. Ultimately the latter leads to a degradation in image quality.
4.4. Watermark Robustness against Attack
Image Processing. We select eight representative types of image-level noise to evaluate the robustness of image processing: (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al., 2020), and (8) Diffusion-based Reconstructive Attack (Zhao et al., 2023d). These cover common image processing techniques, geometric transformations, and regeneration attacks encountered in real-world scenarios. The main results are shown in Table 2.
[…]
PIQE↓ NIQE↓ CLIP↑ FID↓ PIQE↓ NIQE↓ CLIP↑ FID↓
MODEL MISMATCH v2.1 v1.4 63.18 6.87 0.287 87.21 54.7 6.79 0.272 84.53
v1.4 v2.1 59.82 6.72 0.293 72.18 57.8 6.37 0.292 66.28
LORA v2.1 v2.1 53.82 5.72 0.306 69.48 55.82 5.37 0.311 70.98
Textual Inversion v2.1 v2.1 36.87 4.96 0.294 66.47 39.58 4.48 0.285 56.83
Table 4. An experiment on generating images with adversarial watermarks and fine-tuning them with different model versions and methods for personalization. The effectiveness of our adversarial watermark defense through personalized fine-tuning is observed across different model versions and personalized methods.
Main Results. For each image processing attack, we report the average bit accuracy/ TPR@1e-5FPR in Table 2, which shows ConceptWM is indeed robust across all the transformations with the bit accuracy consistently above 0.9. ConceptWM remarkably outperforms existing multi-bit watermarks on the average performance with more than 6.3% boost compared with the state-of-the-art (SoTA) results. Since our method embeds the watermark into the entire latent space and trains the distortion layer, it has a clear advantage over contrastive watermarking schemes when facing various image processing attacks.
PSNR SSIM
BitAcc
BitAcc
GaussianNoise 33.26 0.928 0.962 0.934
ColorJitter 33.78 0.939 0.979 0.972
GaussianBlur 33.69 0.931 0.967 0.948
Jpeg 32.26 0.933 0.969 0.882
CROP&SCALE 30.58 0.917 0.959 0.593
Random Mask 32.19 0.915 0.969 0.908
Combination 29.08 0.889 0.954 0.528
Table 5. The impact of various attacks on image quality in the distortion layer in the watermark components pretraining.
4.5. Concept Protection Evaluation
Model Mismatching. We provide an example of transferring adversarial watermark trained on SD v1.4 and SD v2.1 to defend DreamBooth models trained from v2.1 and v1.4 in Table 4. The results show that our method still works effectively across different versions of the Diffusion Model.
Textual Inversion and DreamBooth with LoRA. Textual Inversion teaches new concepts by adjusting a word vector, rather than fine-tuning the entire model. LoRA, on the other hand, employs low-rank weight updates to enhance memory efficiency and is widely adopted within the community. The relevant results are shown in Table 4. Our method shows better performance in mitigating the degradation effect on generation with LoRA. Since Textual Inversion does not inject concepts directly into the model, the reduction in quality is not significant and the semantic quality of the generation decrease.
Main Results. The results in Table 4 demonstrate that adversarial properties of our watermark is preserved across different versions of Stable Diffusion. When the model is mismatched, PIQE exceeds 57.8 and NIQE exceeds 6.37, indicating that our adversarial properties are inheritable across different versions of the diffusion model. Our method ensures that images generated with LoRA have a PIQE higher than 53.82, while the semantic quality metrics of images generated with Textual Inversion, measured by the CLIP score and FID, are 66.47 and 0.292. It effectively resists other types of personalization fine-tuning attacks, such as LoRA and Textual Inversion.
4.6. Ablation Studies.
Prompt Mismatching. Considering that users generate relevant images based on various prompts combined with specific concepts, we generate images for specific concepts under different prompt conditions. The main results are shown in Table 4, which shows that our watermarking method achieves higher bit extraction accuracy for images similar to the training set. When the image deviates significantly from the set instance prompt, our watermarking method still maintains good precision. Furthermore, the generated watermark exhibits a certain level of adversarial robustness, as the quality of images trained under this watermark is reduced, which is beneficial for the protection of specific concepts in diffusion models.
Distortion Layer. We evaluated the impact of various attacks on image quality in the Distortion Layer in the watermark components pretraining. The main results are in Table 5, which indicates that cropping and the mask adversarial layers have a significant impact on image quality, as they require a stronger watermark to be stably embedded within a smaller space.
Inference Steps. Users may influence watermark extraction under different inference parameters. We generated images of specific concepts at different inference steps and performed watermark extraction. The results, shown in Table 3, indicate that the ConceptWM remains effective even with varying inference steps.
Guidance Scales. Larger guidance scales result in more faithful of the generated image adherence to prompts. We evaluate the impace of guidance scale in the Table 3. The results show that our method maintains good stability across various values of the Guidance scale.
Sampling Method. We evaluate our watermark on DDIM (Song et al., 2020), DPM-Solver multi-step (Lu et al., 2022), Euler and Heun samplers (Karras et al., 2022). The results show that our watermark is minimally affected by the choice of sampler.
Different VAE Decoder. An adversary may interfere with watermark decoding by replacing the VAE decoder. To evaluate the impact of the generated images on the watermark, we used different types of VAEs including stable diffusion V1.4 and V2.1 for decoding. The results in Table 3 indicate that changing the VAE does not have a significant impact on the watermark extraction, which is a clear advantage over VAE-based schemes, as the watermark cannot be preserved when the VAE is replaced in VAE-based methods.
Generation Size. The results in Table 3 show that different generation sizes have a noticeable impact on watermark, while the watermarking method still remains effective and detectable across different generation sizes.
5. Conclusion
This paper reveals the potential threats posed by diffusion model personalization techniques and proposes an adversarial watermarking method to protect specific concepts. Our solution integrates adversarial perturbations and watermarks into the diffusion model, ensuring that images generated by the diffusion model for specific concepts are traceable, and any personalized model trained on these images will produce low-quality results. We have evaluated the stability and adversarial effectiveness of the watermarking method we designed, demonstrating that our defense remains effective under adverse conditions. In the future, our main focus will be on two aspects of improvement. On the one hand, we aim to address the efficiency issue caused by the need to generate adversarial noise in the current method. On the other hand, we will consider potential threats from attackers who may target the model level.