### v3

# Watermarking Visual Concepts for Diffusion Models

Abstract
The personalization techniques of diffusion models succeed in generating images with specific concepts. This ability also poses great threats to copyright protection and network security since malicious users can generate unauthorized content and disinformation relevant to a target concept. Model watermarking is an effective solution to trace the malicious generated images and safeguard their copyright. However, existing model watermarking techniques merely achieve image-level tracing without concept traceability. When tracing infringing or harmful concepts, current approaches execute image concept detection and model tracing sequentially, where performance is critically constrained by concept detection accuracy. In this paper, we propose a lightweight concept watermarking framework that efficiently binds target concepts to model watermarks, supporting simultaneous concept identification and model tracing via single-stage watermark verification. To further enhance the robustness of concept watermarking, we propose an adversarial perturbation injection method collaboratively embedded with watermarks during image generation, avoiding watermark removal by model purification attacks. Experimental results demonstrate that ConceptWM significantly outperforms state-of-the-art watermarking methods, improving detection accuracy by 6.3%-19.3% across diverse datasets including COCO and StableDiffusionDB. Additionally, ConceptWM possesses a critical capability absent in other watermarking methods: it sustains a 21.7% FID/CLIP degradation under adversarial fine-tuning of Stable Diffusion models on WikiArt and CelebA-HQ, demonstrating its capability to mitigate model misuse.
Codehttps://anonymous.4open.science/r/Conceptwm
Introduction
Recent advance in diffusion models have significantly improved image generation ability, enabling the creation of high-quality images from personalized textual descriptions. The upgraded diffusion models achieve few-shot learning for new visual concept generation, such as unique visual themes, motifs, stylistic preferences, or celibrities, using only a few images with the target concepts (Shi et al. 2024; Gal et al. 2023, 2022; Kumari et al. 2023). However, these techniques pose significant risks for visual concept infringement and misuse. For example, a malevolent actor might leverage a few online trademark or digital artwork images to fine-tune off-the-shelf diffusion models and edit the content without permission (Zhu, Takahashi, and Kataoka 2024; Li et al. 2025). Similarly, a few online celebrity images are adequate for malevolent actors to generate misinformation with these celebrities (Wang et al. 2024). It becomes an urgent issue to effectively trace AI-edited visual concept images for copyright protection and privacy preservation.
Figure 1: Five main kinds of diffusion model watermarking methods: (a) Post-processing watermark. (b) VAE finetuning watermark. (c) U-Net finetuning watermark. (d) Latent-based watermark. (e) Concept-oriented watermark.
Model watermarking is a representative way to trace the source model of generated images and thus verify the accountability of malicious content. The typical solutions of model watermarking fine-tune the pre-trained diffusion models to inject watermarks and extrect the watermark information from any generated images by a watermark decoder. Compared with post processing solutions by adding watermarks after image generation (Fig. 1(a)), model watermarking is more secure to avoid watermark evasion (Zhao et al. 2023b). Despite the great progress in diverse model watermarking schemes (Fig. 1(b)(c)(d)), existing schemes indiscriminately inject watermarks into all the generated images agnostic to visual concepts, which has two obvious limitations for visual concept tracing and protection. On the one hand, tracing concept-level malicious image becomes indirect that requires a two-stage pipeline: initial detection of target visual concepts in images, followed by watermark-based tracing for concept images. Thus, the tracing effectiveness is critically constrained by concept detection accuracy. One the other hand, recent studies (Cui et al. 2025) reveal that existing personalization methods enable malicious actors to remove watermarks by fine-tuning diffusion models with a few watermarked samples, while retaining concept generation capabilities. Therefore, current model watermarking methods with only image tracing capability are still unable to prevent misue of watermarked images.
In this paper, we propose an approach that aims to trace and protect generated images with target visual concepts by diffusion model. The core idea of our approach is Concept Watermarking, that is, during image generation, we embed both traceable watermarks and fine-tuning-resistent perturbations into concept-specific generated images. This paradigm enables people to technically govern AI-powered malicous content and proactively defend subsequent misuse of generated content in a finer manner. To this end, three new challenges need to be addressed: First, for a target concept (e.g., ‘a specific kind A of cats’), concept watermarking requires to establish tripartite correlation among a concept watermark (e.g., a bit sequence ‘01100’), visual concepts (diverse appreances of ‘A cats’) in the generated images, and a textual concept (a psuedo word of ‘cat A’) in the prompts, which is quite different from existing watermarking methods that solely establish bipartite correlation between a watermark and model outputs; Second, concept watermarking has adversarial capability that the generated concept images are unable to be used for model fine-tuning, which has no influence on concept-irrelevant generated images; Third, concept watermarking is robust to diverse real-world attacks including image processing, geometric transformations, and model regeneration attacks.
To address these challenges, we propose a lightweight concept watermarking framework, named ConceptWM, which contains two main componets, i.e., efficient concept watermark training and imperceptible adversarial perturbation injection. To embed a watermark that is both fidelity-preserving and conducive to learning by U-Net, we propose a fidelity-preserving latent watermark module that incorporates the watermark into the latent space and introduce a fusion layer after the message encoder and latent concatenation to mitigate rigid watermark pattern artifacts. To achieve concept-level image tracing, we propose an efficient concept watermark training strategy that alternately optimizes DreamBooth fine-tuning and watermark embedding training to achieve joint concept-watermark learning. Our method learns watermark embeddings efficiently, requiring only a few reference images and low training budgets. To prevent the misuse of watermarked images, we propose an imperceptible adversarial perturbation injection module to refine both referenced model parameters and adversarial perturbations by proposed optimization strategy, which balances adversarial robustness, watermark fidelity, and image quality, while mitigating the attack risk by personalized fine-tuning.
Our contributions are as follows: (1) We propose a lightweight concept watermarking framework that imperceptibly embeds watermarks into the generated concept-specific images by diffusion models. Concept watermarking first tackles the challenge of establishing tripartite correlation across three modalities, e.g., watermarks, visual concept and textual concept, which makes it possible for concept-level generated image tracing in one-stop manner. (2) We propose a new adversarial mechanism that endows concept watermarks with proactive denfence capabilities. This mechanism enables the adversarial perturbation injection without interference to image quality and watermark fidelity, while making the generated concept images resistent to malicious use for model fine-tuning. (3) ConceptWM achieves 6.3% higher robustness against image processing than 4 baselines while maintaining a critical property: it forces 21.7% FID/CLIP degradation in fine-tuned Stable Diffusion models - a capability beyond current methods.
Figure 2: Framework of ConceptWM. The ECWT module employs a two-step training process to learn specific concepts and watermark patterns. The IAPI module utilizes a surrogate model fully fine-tuned on a referenced clean image to guide PGD optimization, iteratively generating adversarial watermarks through multiple rounds of watermark and noise integration.
Related Work
Personalization Techniques for Diffusion Models. Recent explorations mainly focus on injecting specific concepts into the model through fine-tunning tools, e.g., Textual Inversion (Gal et al. 2022), DreamBooth (Ruiz et al. 2023) and LoRA (Low-Rank Adaptation) (Hu et al. 2021). Specifically, Textual Inversion focuses on training text embedding so that concepts are allowed to be injected into the text encoder without altering weights of the U-Net. DreamBooth fine-tunes the entire U-Net portion of the Stable Diffusion model and employs prior loss during training to prevent overfitting. LoRA is designed to achieve lightweigth efficient fine-tuning by training incremental weights in the attention layers of the U-Net. Our work focuses on watermarking specific concepts generated through DreamBooth fine-tuning.
Model Watermarking for Latent Diffusion Models. The embedding characteristics of different watermarking methods for diffusion models are illustrated in Figure 1, These watermarking methods aim to protect the copyright of images or generative models, so they often embed watermarks into all images. WDM (Zhao et al. 2023c) is designed to train an autoencoder to stamp a watermark on all training data before re-training the generator from scratch. Aqualora (Feng et al. 2024) is a lora-based method that watermarks all images. VAE-based methods (Xiong et al. 2023; Kim et al. 2024; Fernandez et al. 2023) fine-tune VAE-Decoder to ensure all generated images containing the watermark. However, these methods fail to protect specific concepts. Our Efficient Concept Watermark Training embeds watermark into specific concept with a few images of specific concept.
Protective Perturbation against Stable Diffusion. Recent studies (Wu et al. 2023; Ye et al. 2023; Zhao et al. 2023d; Zheng et al. 2023) mostly have addressed disrupting the fine-tuning process by adding imperceptible protective noise to images, in order to prevent personal images from potential infringement during the fine-tuning of Stable Diffusion. AdvDM employs adversarial noise to safeguard personal images on Stable Diffusion by maximizing the Mean Squared Error loss during optimization (Liang et al. 2023). Anti-DreamBooth (Van Le et al. 2023) offers a bi-level min-max optimization for generating protective perturbations, through integrating the DreamBooth fine-tuning process of Stable Diffusion into its framework. Our Imperceptible Adversarial Perturbation Injection effectively endows the watermark with the aforementioned adversarial properties.
Methodology
The main objective of our scheme is to protect images of a specific concept that are generated by the diffusion model, by embedding an adversarial watermark pattern into the U-Net component of the diffusion model. Fig. 2 illustrates an overview of our approach. To embed a watermark that is both fidelity-preserving and conducive to learning by U-Net, we develop a Fidelity-preserving Latent Watermarking (FLW) module, which adds a combination layer for the message encoder and latent concatenation, so that watermarks can be generated based on image characteristics rather than a fixed watermark signal. To protect the watermarked concept-specific images from misusing, we develop an Imperceptible Adversarial Perturbation Injection module to integrate adversarial propeties into the watermark. To enable U-Net to efficiently learn watermark patterns during concept learning with limited samples, we present an Efficient Concept Watermark Finetuning module that alternates optimization of model parameters to simultaneously achieve adversarial watermark embedding and concept learning.
Fidelity-preserving Latent Watermark
Existing image watermarking techniques frequently arise substantial discrepancies between the watermark embedding pattern and the latent space. This mismatch issue may cause the disruption or even complete loss of watermark information when the image is encoded into the latent space by a VAE encoder or regenerated through diffusion, as has been discussed by recent watermarking attacks (An et al. 2024; Ballé et al. 2018; Cheng et al. 2020). Thus, we aim to embed a robust watermark within the image that is both resilient and conducive to efficient learning for U-Net. To be specific, we have developed two variants of latent watermarking, which are Overlay-based Latent Watermarking (OLW) and Fidelity-preserving Latent Watermarking (FLW). Both variants involve the latent domain perturbations. The watermarked images are processed through the message decoder to evaluate the discrepancy between the decoded watermark information and the original content of the watermarked image. The message loss is the Binary Cross Entropy (BCE) between the message $m$ and $\sigma(m^{\prime})$. Given the message length N, the message loss $\mathcal{L}_{m}$ is defined in Eq. (1).
$\mathcal{L}_{m}=-\sum_{k=1}^{N}m_{k}\log\sigma(m_{k}^{\prime})+(1-m_{k})\log(1-\sigma(m_{k}^{\prime}))$ (1)
Overlay-based Latent Watermarking. OLW essentially overlays a fixed watermark signal that can be removed through subtraction. We observe that OLW-watermarked images may face the issue of the watermark integrity, since noticeable artifacts appear in local regions of the images, making the images distinguishable to the huaman eye.
Fidelity-preserving Latent Watermarking. To address issues of OLW, we have developed Fidelity-preserving Latent Watermarking (FLW) that adds a combination layer after the message encoder and latent concatenation. To ensure visual consistency, we calculate LPIPS loss $\mathcal{L}_{pips}$ (Zhang et al. 2018) and utilize a Peak Pixel Difference (PPD) Loss $\mathcal{L}_{ppd}$ which selects the most prominent parts of the image and calculates channel loss to reduce excessive modification of the images. Our training objective ($\mathcal{L}_{total}$) can be expressed in Eq. (2), where $\lambda$ and $\mu$ represent the coefficients.
$\mathcal{L}_{total}=\mathcal{L}_{m}+\lambda\mathcal{L}_{pips}+\mu\mathcal{L}_{ppd}$ (2)
The jointly optimized message encoder and decoder, working in concert with the Adcersarial Perturbation Modulation module, will empower the U-Net to learn robust watermarking patterns. We provide the training structure comparison of OLW and FLW in the Appendix.
Imperceptible Adversarial Perturbation Injection
To protect the generated concept-specific images from removing through personalized fine-tuning, we integrate adversarial properties into the watermark. By maximizing the image generation loss with adversarial noise added to the watermark and minimizing the concept learning loss through Dreambooth fine-tuning with reference images, we ensure that the watermark does not interfere with concept learning while preventing the misuse of watermarked image.
Image Generation Loss. To enable the conditional denoising model $\epsilon_{\theta}(x,t,c)$ to match the data distribution of the target images, $\mathcal{L}_{cond}$ is a loss function that ensures the conditional denoising model generates images conditioned on a given text condition ($c$). Given the input image $x_{0}$, the $\mathcal{L}_{cond}$ is as follows in Eq. (3):
$\mathcal{L}_{cond}(\theta,x_{0})=\mathbb{E}_{x_{0},t,c,\epsilon\in\mathcal{N}(0,1)}\|\epsilon-\epsilon_{\theta}(x_{t+1},t,c)\|_{2}^{2}$
where each variable $x_{t}$ is constructed through noise injection at a corresponding timestep $t$ within the set $\{1,2,\cdots,T\}$.
[…]
$\mathcal{L}_{ft}(\theta,x_{0})=\mathbb{E}_{x_{0},t}\|\epsilon-\epsilon_{\theta}(x_{t+1},t,c)\|_{2}^{2}+\lambda\|\epsilon^{\prime}-\epsilon_{\theta}(x_{t+1}^{\prime},t,c_{cls})\|_{2}^{2}$
where $\epsilon$ and $\epsilon^{\prime}$ are both sampled from the standard normal distribution $\mathcal{N}(0,\mathbf{I})$, The noisy variable $x_{t+1}^{\prime}$ is generated from the original stable diffusion model based on the prior prompt $c_{\text{cls}}$. $\lambda$ emphasizes the importance of the prior term in the generation process. The term $\lambda$ emphasizes the importance of the prior term in the generation process.
Adversarial Perturbation Injection. The goal of this modulation is to maximize $L_{cond}$ in Eq. (3) to hinder personalized finetuning and minimize $L_{ft}$ in Eq. (4) to learn the concept. Our solution is to use watermark-free images $\mathcal{X}_{A}$ for one-step DreamBooth training to reduce the DreamBooth loss $\mathcal{L}_{\text{ft}}$, while simultaneously applying Projected Gradient Descent (PGD) (Madry 2017) to enhance the adversarial loss $L_{cond}$. The model used for PGD is trained with clean referenced images $\mathcal{X}_{B}$ in a one-step dreambooth training process, rather than the images used for concept learning. The purpose of this approach is to allow the model to learn adversarial perturbations without affecting the learning of the concept. The specific procedure is defined by Eq.s (5) and (6), where $\sigma_{1}$ denotes the residual of the watermarked image and original image. $x^{i}$ denotes the clean samples.
$\theta^{*}=\arg\min_{\theta}\sum_{x\in\mathcal{X}_{A}}\mathcal{L}_{ft}(\theta,x^{i})$ (5)
$\varphi^{(i)}=\arg\max_{\sigma_{2}^{(i)}}\mathcal{L}_{cond}(\theta^{*},x^{(i)}+\sigma_{1}+\sigma_{2}^{(i)})$ (6)
The adversarial noise $\sigma^{*(i)}$ is within the range of $\eta$. The Imperceptible Adversarial Perturbation Modulation outputs an adversarial watermark $\varphi$ as the parameter for Concept-oriented Watermark Training. The pseudo-code of Adversarial Perturbation Modulation is presented in the Appendix.
Efficient Concept Watermark Training
We aim to integrate adversarial watermarking into concept learning in this module. The primary challenges are as follows: The availability of fine-tuned images for specific concepts is often limited, hindering the model’s ability to learn the distinct watermark patterns. The introduction of adversarial perturbations can interfere with the DreamBooth training process, as these protective perturbations prevent the model from effectively learning the concept. Our core idea is to decouple concept learning from adversarial watermark learning. In the DreamBooth stage, the U-Net learns the specific concept. In the adversarial watermark learning stage, the model is fed with both watermarked and non-watermarked images, and by using the difference in the predictions as a loss, it guides the generative model in learning the watermark embedding pattern.
The training objective of this stage is defined by Eq. (7), where $\varphi$ represents the adversarial perturbation containing the watermark generated through IAPI, $\epsilon_{\theta}^{*}$ denotes the fixed model, and $\epsilon_{\theta}$ refers to the trained model.
$\mathcal{L}_{\mathrm{cwft}}(\theta,x_{t})$ $=\mathbb{E}_{t,c,x}\Big{[}\|\boldsymbol{\epsilon}_{\theta}(x_{t}+\varphi,t,c)-\boldsymbol{\epsilon}_{\theta}^{*}(x_{t},t,c\big{)}\|^{2}\Big{]}$ (7)
The pseudo-code of Concept-oriented Watermarking Training is presented in the Appendix.
[…]
Figure 3: Demonstration of the generated images. Below each image, we display the corresponding Bit Accuracy, along with the image quality metrics PIQE and NIQE.
Type Method
Concept
Tracing
Concept
Protection
Fidelity Robustness
FID CLIP BitAcc BitAcc(Adv.) TPR TPR (Adv.)
None - - 24.37 0.367 - - - -
Post-Processing DwtDctSvd $\times$ $\times$ 25.01 0.363 99.98 71.47 1.000 0.367
RivaGan $\times$ $\times$ 24.72 0.363 98.47 83.21 0.994 0.651
Latent-based Tree-Ring $\times$ $\times$ 25.69 0.364 - - 1.000 0.821
VAE-based StableSig. $\times$ $\times$ 24.92 0.362 98.49 81.94 0.993 0.848
Concept-Oriented OLW-based 24.06 0.357 98.68 90.03 0.998 0.884
FLW-based 24.58 0.365 95.68 89.32 0.994 0.858
Table 1: A comparison between our method and previous watermarking techniques is presented. We control the false positive rate (FPR) at $10^{-5}$ and evaluate the true positive rate (TPR). Since Tree-ring is a zero-bit watermark, bit accuracy cannot be calculated in this case. “Adv.” (Adversarial) refers to the average performance of images under various distortions.
Experiment
Datasets. To evaluate the effectiveness of the proposed methods, for subject-driven generation, we select the face dataset CelebA-HQ (Zhu et al. 2022). For style-driven generation, we select 12 artists with distinct styles from WikiArt  [^1] https://www.wikiart.org, with each artist contributing 10 artworks that are consistent in terms of both artistic style and time period. Each image is resized and center-cropped to a resolution of 512 × 512. In each experiment, a total of 5–8 carefully selected images were utilized for DreamBooth fine-tuning.
[…]
Metrics Methods Distortions
Contrast Brightness Blur Noise Jpeg C.&R. Vae.C Denoising Average
Bit Accuracy(%) Dwtdctsvd 88.41 86.80 92.42 65.87 83.32 54.02 55.74 49.53 72.01
RivaGan 97.96 94.67 94.89 91.09 90.68 94.17 70.62 58.64 86.59
Stablesig. 97.86 95.48 89.88 78.94 84.40 97.47 69.94 52.58 83.31
Ours 94.26 93.46 93.68 92.42 93.63 90.26 90.12 84.63 91.34
[…]
RivaGan 0.989 0.956 0.979 0.926 0.928 0.956 0.284 0.010 0.754
Stablesig. 0.987 0.985 0.905 0.369 0.821 0.989 0.293 0.000 0.669
Tree-ring 1.000 1.000 0.954 0.701 0.905 0.121 0.958 1.000 0.830
Ours 0.968 0.951 0.947 0.958 0.949 0.912 0.910 0.734 0.916
Table 2: The performance of ConceptWM against various types of attacks is compared with other watermarking schemes. The attacks include (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al. 2020) and (8) Diffusion-based Reconstructive Attack (Zhao et al. 2023a).
Evaluation Metrics. We evaluate the detection performance using the true positive rate (TPR) at a false positive rate (FPR) of 10e-5. For traceability performance, we assess bit accuracy. To measure the quality of image generation, we compute the Peak Signal-to-Noise Ratio (PSNR) (Hore and Ziou 2010) and Structural Similarity Index (SSIM) (Wang et al. 2004); and the Natural Image Quality Evaluator (NIQE) (Mittal, Soundararajan, and Bovik 2012) and Perceptual Image Quality Evaluator (PIQE) (Venkatanath et al. 2015) for evaluating overall image quality.
Watermarking baselines. We select four typical baselines: two official watermark of Stable Diffusion (Rombach et al. 2022) called DwtDctSvd (Cox et al. 2007), and RivaGAN (Zhang et al. 2019),VAE-based method Stable Signature (Fernandez et al. 2023) and Latent-based method Tree-ring (Wen et al. 2024). DwtDctSvd, RivaGAN, and Stable Signature all use 48-bit watermarks, while Tree-ring uses P-value as the evaluation method.
Watermark Performance Comparison
We evaluate our method against 4 typical baselines on two tasks: (1) Detection: All methods are tested as single-bit watermarks with a unified watermark, with the False Positive Rate (FPR) set to $10^{-5}$, and we assess the True Positive Rate (TPR) on 1,000 watermarked images. (2) Traceability: Except for the single-bit watermark Tree-Ring, all other methods are tested as multi-bit watermarks. Specifically, we simulate a scenario with 1,000 model users, each requiring a distinct watermark for model tracing. Each user generates 10 images, creating a dataset of 10,000 watermarked images. During testing, we calculate the Bit Accuracy by comparing the watermark of each user with the watermarked images generated by the user.
The comparison results, shown in Table 1, demonstrate that our watermarking method achieves superior robustness, significantly outperforming the baselines in both tasks. In terms of Bit Accuracy, our method outperforms the best baseline by around 6.3%. OLW exhibits strong performance in robustness experiments. However, it suffers from relatively unsatisfactory image quality (FID 24.06,CLIP 0.357). FLW demonstrates a relatively balanced result, maintaining good image quality while ensuring strong robustness. We found that there is a trade-off between image quality, robustness, and detection accuracy in watermark embedding. Due to extensive adversarial training on the watermark model, FLW achieves a bit accuracy of 95.68 in a non-attack environment, which is relatively lower compared to other watermarking schemes. To further improve bit accuracy in a non-attack setting, it would require considering either lower image quality or reduced robustness.
Watermarked Diffusion Model Fidelity
For fidelity evaluation metrics, we use the Fréchet Inception Distance (FID) (Heusel et al. 2017) for assessing image diversity and semantic relevance, which is calculated on 5000 images from the COCO 2017 validation set (Lin et al. 2014). Additionally, we use CLIP score (Radford et al. 2021), a method for measuring image similarity, which provides results more aligned with human judgment. Table 1 presents a comparison of our method with other baseline methods.
Settings Bit Accuracy TPR
Steps 25 0.917 0.921
50 0.936 0.946
100 0.936 0.951
Sampler DDIM 0.929 0.939
DPM-M 0.938 0.972
EULER 0.922 0.941
HEUN 0.947 0.962
Guidance Scale 5 0.932 0.948
7.5 0.917 0.932
10 0.933 0.947
Generation Size 512 0.933 0.951
768 0.812 0.748
VAE V2.1 0.939 0.954
V1.4 0.901 0.913
Table 3: The impact of inference settings on ConceptWM watermark detection.
Main Results. The performance of our method outperforms the state-of-the-art methods, with minimal performance difference compared to the original model (FID difference of 0.21, CLIP difference of 0.02). This method is a localized fine-tuning of the model and does not significantly impact the model’s overall performance. Due to the smaller dataset and the more refined watermarking method, our method has a smaller impact on model generation quality compared to other methods. It can be observed that, among the two variants of latent watermarking introduced, OLW and FLW, the image quality produced by OLW, with an FID of 24.06 and semantic quality (CLIP) of 0.357, is comparatively lower. In contrast, the image quality produced by FLW, has an FID of 24.58 and semantic quality (CLIP) of 0.365. This disparity arises because the watermark in FLW is contextually related to the image content, whereas the watermark in OLW is a fixed, image-independent watermark, which leads to degradation in image quality.
Watermark Robustness against Attack
Image Processing. We select eight representative types of image-level noise to evaluate the robustness of image processing: (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al. 2020), and (8) Diffusion-based Reconstructive Attack (Zhao et al. 2023b). These cover common image processing techniques, geometric transformations, and regeneration attacks encountered in real-world scenarios. The main results are shown in Table 2.
[…]
PIQE↓ NIQE↓ CLIP↑ FID↓ PIQE↓ NIQE↓ CLIP↑ FID↓
MODEL
MISMATCH
v2.1 v1.4 63.18 6.87 0.287 87.21 54.7 6.79 0.272 84.53
v1.4 v2.1 59.82 6.72 0.293 72.18 57.8 6.37 0.292 66.28
LORA v2.1 v2.1 53.82 5.72 0.306 69.48 55.82 5.37 0.311 70.98
Textual Inversion v2.1 v2.1 36.87 4.96 0.294 66.47 39.58 4.48 0.285 56.83
Table 4: An experiment on generating images with adversarial watermarks and fine-tuning them with different model versions and methods for personalization. The effectiveness of our adversarial watermark defense through personalized fine-tuning is observed across different model versions and personalized methods.
Main Results. For each image processing attack, we report the average bit accuracy/ TPR@1e-5FPR in Table 2, which shows ConceptWM is indeed robust across all the transformations with the bit accuracy consistently above 0.9. ConceptWM remarkably outperforms existing multi-bit watermarks on the average performance with more than 6.3% boost compared with the state-of-the-art (SoTA) results. Since our method embeds the watermark into the entire latent space and trains the distortion layer, it has a clear advantage over contrastive watermarking schemes when facing various image processing attacks.
PSNR SSIM
BitAcc
(w/)
BitAcc
(w/o)
GaussianNoise 33.26 0.928 0.962 0.934
ColorJitter 33.78 0.939 0.979 0.972
GaussianBlur 33.69 0.931 0.967 0.948
Jpeg 32.26 0.933 0.969 0.882
CROP&SCALE 30.58 0.917 0.959 0.593
Random Mask 32.19 0.915 0.969 0.908
Combination 29.08 0.889 0.954 0.528
Table 5: The impact of various attacks on image quality in the distortion layer during pretraining.
Concept Protection Evaluation
Model Mismatching. We provide an example of transferring adversarial watermark trained on SD v1.4 and SD v2.1 to defend DreamBooth models trained from v2.1 and v1.4 in Table 4. The results show that our method still works effectively across different versions of the Diffusion Model.
Textual Inversion and DreamBooth with LoRA. Textual Inversion teaches new concepts by learning a new token, rather than fine-tuning the entire model. LoRA, on the other hand, employs low-rank weight updates to enhance memory efficiency. The relevant results are shown in Table 4. Our method shows better performance in mitigating the degradation effect on generation with LoRA. Since Textual Inversion does not inject concepts directly into the model, the reduction in quality is not significant and the semantic quality of the generation decrease.
Main Results. The results in Table 4 demonstrate that adversarial properties of our watermark is preserved across different versions of Stable Diffusion. When the model is mismatched, PIQE exceeds 57.8 and NIQE exceeds 6.37, indicating that our adversarial properties are inheritable across different versions of the diffusion model. Our method ensures that images generated with LoRA have a PIQE higher than 53.82, while the semantic quality metrics of images generated with Textual Inversion, measured by the CLIP score and FID, are 66.47 and 0.292. It effectively resists other types of personalization fine-tuning attacks, such as LoRA and Textual Inversion.
Ablation Studies
Prompt Mismatching. Considering that users generate relevant images based on various prompts combined with specific concepts, we generate images for specific concepts under different prompt conditions. The results are shown in Table 3. When the image deviates significantly from the set instance prompt, our watermarking method still maintains bit accuacy higher than 0.8.
Distortion Layer. We evaluate the impact of Distortion Layer in the watermark components pretraining. The main results are in Table 5, which indicates that cropping and the mask adversarial layers have a significant impact on image quality, as they require a stronger watermark to be stably embedded within a smaller space.
Inference Steps. As shown in Table 3, the bit accuracy remains stable at approximately 0.93 across different settings - specifically reaching 0.917 at 25 steps, 0.936 at 50 steps, and 0.936 at 100 steps. The results indicate that the ConceptWM remains effective with varying inference steps.
Guidance Scales. Our experiments demonstrate that ConceptWM maintains robust watermark detection performance across varying guidance scales, achieving 0.932 bit accuracy at scale 5, 0.917 at scale 7.5, and 0.933 at scale 10.
Sampling Method. Evaluation on DDIM (Song, Meng, and Ermon 2020), DPM-Solver (Lu et al. 2022), Euler and Heun samplers (Karras et al. 2022) shows consistent performance with 0.929, 0.938, 0.922 and 0.947 bit accuracy respectively, demonstrating minimal impact from sampler choice.
Different VAE Decoder. An adversary may interfere with watermark decoding by replacing the VAE decoder. We used different types of VAEs including stable diffusion V1.4 and V2.1 for decoding, achieving 0.901 and 0.939 bit accuracy respectively. The results in Table 3 indicate that changing the VAE does not have a significant impact on the watermark extraction, which is a clear advantage over VAE-based schemes, as the watermark cannot be preserved when the VAE is replaced in VAE-based methods.
Generation Size. The results in Table 3 show that different generation sizes have some impact on watermark accuracy, with 0.933 for 512×512 and 0.812 for 768×768 generations, while the watermarking method still remains effective and detectable across different generation sizes.
Conclusion
This paper reveals the potential threats posed by diffusion model personalization techniques and proposes an adversarial watermarking method to protect specific concepts. Our solution integrates adversarial perturbations and watermarks into the diffusion model, ensuring that images generated by the diffusion model for specific concepts are traceable, and any personalized model trained on these images will produce low-quality results. In the future, The investigation of concept-specific learning can be extended beyond DreamBooth, necessitating an in-depth examination of watermark embedding paradigms for textual inversion, LoRA, and related concept learning methodologies.